# High-Performance Multi-Qubit System with Double-Transmon Couplers towards Scalable Superconducting Quantum Computers

### v1

@@ VI Summary
In this paper, we have analyzed the system of three fixed-frequency-transmon qubits coupled via two DTCs, by numerical simulation using the circuit model, where the NN qubits are highly detuned and NNN qubits are nearly resonant. As a result, we have found that the DTC can suppress the residual $ZZ$ couplings and $ZZZ$ coupling, as well as in isolated two-qubit systems. Moreover, we have succeeded in implementing 30-ns CZ gates and 10-ns $\pi/2$ pulses with average fidelities of 0.9999 or higher. The degradation of the CZ gate fidelities due to the increasing number of qubits from two to three is smaller than 0.0001.
For comparison, additional simulations of the system where DTCs were replaced by STCs have been performed. Then, we have found that both suppressing residual couplings to several kHz and achieving two-qubit and single-qubit gates with fidelities of 0.9999 are highly challenging for the STCs. Due to the structure of STCs, the NNN-qubit coupling strength $g_{13}$ tends to be larger than DTCs, and hence causes more severe errors.
@@ Appendix C Dimension reduction technique
In this section, we explain our numerical method to reduce the computational cost. Here we only explain the case of the three-qubit system with two DTCs. Other systems can also be treated similarly.
@@ C.1 Diagonalization
The size of these matrces is $(N_{123}\times N_{45}\times N_{67})\times(N_{123}\times N_{45}\times N_{67})=75000\times 75000$. By diagonalizing $\hat{H}^{\prime}$, we can obtain the energy eigenvalues of the total system.
@@ C.2 Gate simulation
From the results of the diagonalization of the total Hamiltonian explained in the previous subsection, we identify the idling point. Fixing $\Theta_{\rm ex,R}$ to $\Theta_{\rm Id,R}$, we again diagonalize $\hat{H}^{\prime}$ for $\Theta_{\rm ex,L}=\Theta_{n}=0.05n\pi\ (n=0,1,\cdots,20)$ to obtain the eigenenergy matrix $\hat{e}_{0}(\Theta_{n})$ and the energy eigenstate matrix $\hat{V}_{0}(\Theta_{n})$ satisfiyng the following relation:
$\hat{H}^{\prime}(\Theta_{n})\hat{V}_{0}(\Theta_{n})=\hat{e}_{0}(\Theta_{n})\hat{V}_{0}(\Theta_{n}),$ (80)
@@ D.1 DTC architecture
$\zeta_{12}$ $=\omega^{\rm L}_{1,1}-\left(\omega^{\rm L}_{1,0}+\omega^{\rm L}_{0,1}\right),$ (86)
$\zeta_{23}$ $=\omega^{\rm R}_{1,1}-\left(\omega^{\rm R}_{1,0}+\omega^{\rm R}_{0,1}\right).$ (87)
These values should be as small as possible at the idling point to suppress conditional phase rotation which causes quantum crosstalk. As shown in Fig. 3(c)[3(d)] by the scatter plot, the minimum value of $|\zeta_{12(23)}|/(2\pi)$ is about 2(2)222(2)2 ( 2 ) kHz at $\Theta_{\rm ex,L(R)}=0.65\pi(0.65\pi)\equiv\Theta_{\rm Id,L(R)}$. Hence, we can choose $\Theta_{\rm Id,L}$ and $\Theta_{\rm Id,R}$ as idling points of system L and R, respectively.
@@ D.2 STC architecture
Similarly to the DTC architecture, we define $\hat{H}^{\rm L(R)}$. Its eigenfrequencies and corresponding eigenstates are denoted by $\omega^{\rm L}_{Q_{1},Q_{2},C_{4}}\ \left(\omega^{\rm R}_{Q_{2},Q_{3},C_{5}}\right)$ and |$\ket{Q_{1},Q_{2},C_{4}}_{\rm L}\ \left(\ket{Q_{2},Q_{3},C_{5}}_{\rm R}\right)$, respectively. We also use notatations $\omega^{\rm L}_{Q_{1},Q_{2}}\equiv\omega^{\rm L}_{Q_{1},Q_{2},0}\ \left(\omega^{\rm R}_{Q_{2},Q_{3}}\equiv\omega^{\rm R}_{Q_{2},Q_{3},0}\right)$ and |$\ket{Q_{1},Q_{2}}_{\rm L}\equiv\ket{Q_{1},Q_{2},0}_{\rm L}\ \left(\ket{Q_{2},Q_{3}}_{\rm R}\equiv\ket{Q_{2},Q_{3},0}_{\rm R}\right)$. Eigenfrequencies of the computational states in the system L (R), $\omega^{\rm L(R)}_{0,0},\omega^{\rm L(R)}_{1,0},\omega^{\rm L(R)}_{0,1},\omega^{\rm L(R)}_{1,1}$, as functions of $\Theta_{\rm ex,L(R)}$ are shown by scatter plots in Fig. 5(a)[5(b)] where we have set $\omega_{0,0}^{\rm L(R)}$ to 0 (outside the graph scale).
The $ZZ$ couplings, $\zeta_{12}$ in the system L and $\zeta_{23}$ in the system R are calculated by Eq. (86) and Eq. (87), respectively. As shown in Fig. 5(c)[5(d)] by the scatter plot, the minimum value of $|\zeta_{12(23)}|/(2\pi)$ is about 64(64)646464(64)64 ( 64 ) kHz at $\Theta_{\rm ex,L(R)}=0(0)\equiv\Theta_{\rm Id,L(R)}$. Hence, we choose $\Theta_{\rm Id,L}$ and $\Theta_{\rm Id,R}$ as idling points of system L and R, respectively. The residual couplings are summarized in Table 5. They are in good agreement with the result of Ref. ETH, : about $-60$ kHz. Their magnitudes are roughly 30 times larger than the ones of the corresponding subsystems with the DTC.
@@ E.2 Two-qubit systems
Figure 6: Optimized gate pulses. The left and right column correspond to the DTC and TC architectures, respectively. (a) and (b): optimized $\Theta_{\rm ex,L}(t)$. (c) and (d): optimized $\Theta_{\rm ex,R}(t)$. In (a)–(d), solid curves and scatter plots represent the pulses in three-qubit systems and two-qubit subsystems, respectively. (e) and (f): optimized $\alpha_{1}(t)$. (g) and (h): optimized $\alpha_{2}(t)$. (i) and (j): optimized $\alpha_{3}(t)$.
@@ F.1 CZ gate
$\hat{u}^{\prime}$ $=\begin{pmatrix}1&\ &\ &\ \\ \ &e^{i\phi_{010}}&\ &\ \\ \ &\ &e^{i\phi_{100}}&\ \\ \ &\ &\ &-e^{i(\phi_{010}+\phi_{100})}\\ \end{pmatrix}\otimes\begin{pmatrix}1&0\\ 0&e^{i\phi_{001}}\\ \end{pmatrix}$ (114)
@@ F.2 $\pi/2$ pulse
$\hat{u}^{\prime}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-e^{i\phi_{A}}\\ e^{i(\phi_{100}-\phi_{A})}&e^{i\phi_{100}}\\ \end{pmatrix}\otimes\begin{pmatrix}1&0\\ 0&e^{i\phi_{010}}\\ \end{pmatrix}\otimes\begin{pmatrix}1&0\\ 0&e^{i\phi_{001}}\\ \end{pmatrix}$ (127)
$=\hat{P}_{Z}^{\dagger}(\phi_{100}-\phi_{A},\phi_{010},\phi_{001})\hat{U}_{\pi/2,1}\hat{P}_{Z}(-\phi_{A},0,0),$ (128)
where $\phi_{A}=\arg(-u^{\prime}_{04})$. Similar caliburations are valid for the $\pi/2$ pulses for ${\rm Q_{2}}$ and ${\rm Q_{3}}$.
@@ G.2 $\pi/2$ pulse
It is known that applying a non-zero $\alpha_{i}$ modulated at the qubit frequency $\omega_{{\rm Id},i}$ of ${\rm Q}_{i}$ at the idling point induces a Rabi ocillation in the ${\rm Q}_{i}$ subspace. To supress the nonadiabatic error, we asuume the following DRAG-like pulse shapeDRAG :
@@
where $\eta_{i}$ is the anharmonicity of the ${\rm Q}_{i}$ at the idling point, and $T$ is the gate time. The values $a_{i}$ and $b_{i}$ are hyperparameters that determine the pulse shape. Applying the above $\alpha_{i}$, we calculate infidelity 1-$\bar{F}_{\pi/2,i}$ as a cost function, and optimize $a_{i}$ and $b_{i}$ by using python version of the L-BFGS-B optimizer of the optimparallel packageoptimparallel1 ; optimparallel2 . We obtain the optimal pulse as the one with the lowest infidelity [see Figs. 6(e)-6(j) for their shapes].
@@ Appendix H Error analysis
Figure 7: Qubit frequencies during the CZ gate. (a) and (b) show qubit frequencies, $\omega_{1,0,0}$ and $\omega_{0,0,1}$, as functions of $\Theta_{\rm ex,L}$ and $\Theta_{\rm ex,R}$, respectively. The inset figures show typical flux-pulse shapes [$\Theta_{\rm ex,L}(t)$ and $\Theta_{\rm ex,R}(t)$] for implementing CZ gates. The vertical dotted lines indicate the pulse peaks. The gray-solid lines and gray-dashed ones with arrows indicate trajectories of $\omega_{1,0,0}$ and $\omega_{0,0,1}$, respectively. Their trajectories are crossed only in (b). Note that $\Theta_{\rm ex,R}$ in (a) and $\Theta_{\rm ex,L}$ in (b) are fixed to the idling points.
@@ H.1 Microwave crosstalk during single-qubit gate operation
We find that, in the DTC architecture, $|\theta_{13}|\sim|\theta_{31}|\sim 10^{-3}\pi$ and the other rotation angles are equal to or smaller than $10^{-5}\pi$. They are small enough to implement $\pi/2$ pulses with a fidelity over 0.9999.
@@
As a result, we find that all the off-diagonal elements of $\hat{U}^{\prime}$ have an absolute value smaller than $10^{-2}$ in the DTC architecture. In other words, undesired transitions are almost negligible during a CZ gate in the DTC architecture, leading to the average gate fidelity of 0.9999.

### v2

@@ VI Summary
In this paper, we have analyzed the system of three fixed-frequency-transmon qubits coupled via two DTCs, by numerical simulation using the circuit model, where the NN qubits are highly detuned and NNN qubits are nearly resonant. As a result, we have found that the DTC can suppress the residual $ZZ$ couplings and $ZZZ$ coupling, as well as in isolated two-qubit systems. Moreover, we have succeeded in implementing 30-ns CZ gates and not only individual but also simultaneous 10-ns $\pi/2$ pulses with average fidelities over 99.99%. The degradation of the CZ gate fidelities due to the increasing number of qubits from two to three is negligible in the 4-digit precision.
For comparison, additional simulations of the system where DTCs were replaced by STCs have been performed. Then, we have found that both suppressing residual couplings to several kHz and achieving two-qubit and single-qubit gates with fidelities of 99.99% are highly challenging for the STCs. Due to the structure of STCs, the NNN-qubit coupling strength $g_{13}$ tends to be larger than DTCs, and hence causes more severe errors.
@@ Appendix C Dimension reduction technique
In this section, we introduce our numerical method to reduce the computational cost. Here we only explain the case of the three-qubit system with two DTCs. Other systems can also be treated similarly.
@@ C.1 Diagonalization
The size of these matrices is $(N_{123}\times N_{45}\times N_{67})\times(N_{123}\times N_{45}\times N_{67})=75000\times 75000$. By diagonalizing $\hat{H}^{\prime}$, we can obtain the energy eigenvalues of the total system.
@@ C.2 Gate simulation
From the results of the diagonalization of the total Hamiltonian explained in the previous subsection, we identify the idle point. Fixing $\Theta_{\rm ex,R}$ to $\Theta_{\rm Id,R}$, we again diagonalize $\hat{H}^{\prime}$ for $\Theta_{\rm ex,L}=\Theta_{n}=0.05n\pi\ (n=0,1,\cdots,20)$ to obtain the eigenenergy matrix $\hat{e}_{0}(\Theta_{n})$ and the energy eigenstate matrix $\hat{V}_{0}(\Theta_{n})$ satisfiyng the following relation:
$\hat{V}_{0}(\Theta_{n})^{\dagger}\hat{H}^{\prime}(\Theta_{n})\hat{V}_{0}(\Theta_{n})=\hat{e}_{0}(\Theta_{n}),$ (67)
@@ D.1 DTC architecture
$\zeta_{12}$ $=\omega^{\rm L}_{1,1}-\omega^{\rm L}_{1,0}-\omega^{\rm L}_{0,1}+\omega^{\rm L}_{0,0},$ (73)
$\zeta_{23}$ $=\omega^{\rm R}_{1,1}-\omega^{\rm R}_{1,0}-\omega^{\rm R}_{0,1}+\omega^{\rm R}_{0,0}.$ (74)
These values should be as small as possible at the idle point to suppress conditional phase rotation which causes quantum crosstalk. As shown in Fig. 3(c)[3(d)] by the scatter plot, the minimum value of $|\zeta_{12(23)}|/(2\pi)$ is about 2(2)222(2)2 ( 2 ) kHz at $\Theta_{\rm ex,L(R)}=0.65\pi(0.65\pi)\equiv\Theta_{\rm Id,L(R)}$. Hence, we can choose $\Theta_{\rm Id,L}$ and $\Theta_{\rm Id,R}$ as idle points of system L and R, respectively.
@@ D.2 STC architecture
Similarly to the DTC architecture, we define $\hat{H}^{\rm L(R)}$. Its eigenfrequencies and corresponding eigenstates are denoted by $\omega^{\rm L}_{Q_{1},Q_{2},C_{4}}\ \left(\omega^{\rm R}_{Q_{2},Q_{3},C_{5}}\right)$ and |$\ket{Q_{1},Q_{2},C_{4}}_{\rm L}\ \left(\ket{Q_{2},Q_{3},C_{5}}_{\rm R}\right)$, respectively. We also use notatations $\omega^{\rm L}_{Q_{1},Q_{2}}\equiv\omega^{\rm L}_{Q_{1},Q_{2},0}\ \left(\omega^{\rm R}_{Q_{2},Q_{3}}\equiv\omega^{\rm R}_{Q_{2},Q_{3},0}\right)$ and |$\ket{Q_{1},Q_{2}}_{\rm L}\equiv\ket{Q_{1},Q_{2},0}_{\rm L}\ \left(\ket{Q_{2},Q_{3}}_{\rm R}\equiv\ket{Q_{2},Q_{3},0}_{\rm R}\right)$. Eigenfrequencies of the computational states in the system L (R), $\omega^{\rm L(R)}_{0,0},\omega^{\rm L(R)}_{1,0},\omega^{\rm L(R)}_{0,1},\omega^{\rm L(R)}_{1,1}$, as functions of $\Theta_{\rm ex,L(R)}$ are shown by scatter plots in Fig. 6(a)[6(b)] where we have set $\omega_{0,0}^{\rm L(R)}$ to 0 (outside the graph scale).
The $ZZ$ couplings, $\zeta_{12}$ in the system L and $\zeta_{23}$ in the system R are calculated by Eq. (73) and Eq. (74), respectively. As shown in Fig. 6(c)[6(d)] by the scatter plot, the minimum value of $|\zeta_{12(23)}|/(2\pi)$ is about 64(64)646464(64)64 ( 64 ) kHz at $\Theta_{\rm ex,L(R)}=0(0)\equiv\Theta_{\rm Id,L(R)}$. Hence, we choose $\Theta_{\rm Id,L}$ and $\Theta_{\rm Id,R}$ as idle points of system L and R, respectively. The residual couplings are summarized in Table 5. They are in good agreement with the result of Ref. ETH, : about $-60$ kHz. Their magnitudes are roughly 30 times larger than the ones of the corresponding subsystems with the DTC.
@@ E.2 Two-qubit systems
Figure 7: Optimized gate pulses. The left and right column correspond to the DTC and STC architectures, respectively. (a) and (b): optimized $\Theta_{\rm ex,L}(t)$. (c) and (d): optimized $\Theta_{\rm ex,R}(t)$. In (a)–(d), solid curves and scatter plots represent the pulses in three-qubit systems and two-qubit subsystems, respectively. (e) and (f): optimized $\alpha_{1}(t)$. (g) and (h): optimized $\alpha_{2}(t)$. (i) and (j): optimized $\alpha_{3}(t)$.
@@ F.1 CZ gate
$\hat{u}^{\prime}$ $=\begin{pmatrix}1&\ &\ &\ \\ \ &e^{i\phi_{010}}&\ &\ \\ \ &\ &e^{i\phi_{100}}&\ \\ \ &\ &\ &-e^{i(\phi_{010}+\phi_{100})}\\ \end{pmatrix}\otimes\begin{pmatrix}1&0\\ 0&e^{i\phi_{001}}\\ \end{pmatrix}$
@@ F.2 $\pi/2$ pulse
$\hat{u}^{\prime}$ $=\frac{1}{\sqrt{2}}\begin{pmatrix}1&-ie^{i\phi_{A}}\\ -ie^{i(\phi_{100}-\phi_{A})}&e^{i\phi_{100}}\\ \end{pmatrix}\otimes\begin{pmatrix}1&0\\ 0&e^{i\phi_{010}}\\ \end{pmatrix}\otimes\begin{pmatrix}1&0\\ 0&e^{i\phi_{001}}\\ \end{pmatrix}$
$=\hat{P}_{Z}^{\dagger}(\phi_{100}-\phi_{A},\phi_{010},\phi_{001})\hat{U}_{\pi/2,\{1\}}\hat{P}_{Z}(-\phi_{A},0,0),$ (91)
where $\phi_{A}=\arg(iu^{\prime}_{04})$. Similar caliburations are valid for the $\pi/2$ pulses for ${\rm Q_{2}}$ and ${\rm Q_{3}}$.
@@ G.2 $\pi/2$ pulse
It is known that applying a non-zero $\alpha_{i}$ modulated at the qubit frequency $\omega_{{\rm Id},i}$ of ${\rm Q}_{i}$ at the idle point induces a Rabi ocillation in the ${\rm Q}_{i}$ subspace. To supress the nonadiabatic error, we asuume the following DRAG-like pulse shapeDRAG :
@@
where $\eta_{i}$ is the anharmonicity of the ${\rm Q}_{i}$ at the idle point, and $T$ is the gate time. The values $a_{i}$ and $b_{i}$ are hyperparameters that determine the pulse shape. Applying the above $\alpha_{i}$, we calculate infidelity 1-$\bar{F}_{\pi/2,\{i\}}$ as a cost function, and optimize $a_{i}$ and $b_{i}$ by using python version of the L-BFGS-B optimizer of the optimparallel packageoptimparallel1 ; optimparallel2 . We obtain the optimal pulse as the one with the lowest infidelity [see Figs. 7(e)-7(j) for their shapes].
@@ Appendix H Error analysis
Figure 8: Qubit frequencies during the CZ gate. (a) and (b) show qubit frequencies, $\omega_{1,0,0}$ and $\omega_{0,0,1}$, as functions of $\Theta_{\rm ex,L}$ and $\Theta_{\rm ex,R}$, respectively. The inset figures show typical flux-pulse shapes [$\Theta_{\rm ex,L}(t)$ and $\Theta_{\rm ex,R}(t)$] for implementing CZ gates. The vertical dotted lines indicate the pulse peaks. The gray-solid lines and gray-dashed ones with arrows indicate trajectories of $\omega_{1,0,0}$ and $\omega_{0,0,1}$, respectively. Their trajectories are crossed only in (b). Note that $\Theta_{\rm ex,R}$ in (a) and $\Theta_{\rm ex,L}$ in (b) are fixed to the idle points.
@@ H.1 Microwave crosstalk during single-qubit gate operation
We find that, in the DTC architecture, $|\theta_{13}|\sim|\theta_{31}|\sim 10^{-3}\pi$ and the other rotation angles are equal to or smaller than $10^{-5}\pi$. They are small enough to implement $\pi/2$ pulses with a fidelity over 99.99%.
@@
As a result, we find that all the off-diagonal elements of $\hat{U}^{\prime}$ have an absolute value smaller than $10^{-2}$ in the DTC architecture. In other words, undesired transitions are almost negligible during a CZ gate in the DTC architecture, leading to the average gate fidelity of 99.99%.
@@
Figure 9: (a) A three-qubit system of diagonal structure (highlighted part). (b)-(e) $ZZ$ and $ZZZ$ couplings around the idle point as functions of $\Theta_{\rm ex,L}.$ Here, $\Theta_{\rm ex,R}$ is fixed to $\Theta_{\rm Id,R}$. (f) The average gate fidelities of 30-ns CZ gates and 10-ns $\pi/2$ pulses, implemented by optimized pulses, as functions of $\kappa$.
Appendix I Diagonal setting
In the main text, we considered the three-qubit systems in which the NNN qubits are located on the horizontal line like Fig. 1(a). To scale up a superconducting quantum computer, the diagonal structure such as the highlighted part of Fig. 9(a) is also important. Since the distances $\rm{Q_{1}}$-$\rm{Q_{3}}$, $\rm{Q_{1(3)}}$-DTC R(L), and DTC L-DTC R in this structure tend to be closer than the ones of the horizontal structure, the parasitic capacitances between them tend to increase. Therefore, it is necessary to clarify how the performance of the three-qubit system changes in response to the increases of the parasitic capacitances. Here, we investigate the case where the above parasitic capacitances, $C_{ij}=C_{ji}\ {\rm for}\ ij\in\{13,16,17,34,35,46,47,56,57\}$, are increased by a factor of $\kappa\in{[1.0,3.0]}$ as follows:
$C_{ij(ji)}\rightarrow\kappa C_{ij(ji)}.$ (96)
Figures 9(b)-9(e) show the $ZZ$ and $ZZZ$ couplings around the idle point. We find that the residual couplings do not increase by more than 1 kHz up to $\kappa=3$. It means that the ability of the residual coupling suppression of the DTC is highly robust against the increase of the parasitic capacitances.
Figure 9(f) shows the average gate fidelities of 30-ns CZ gates and 10-ns $\pi/2$ pulses, implemented by optimized pulses, as functions of $\kappa$. The fidelities tend to decrease as $\kappa$ increases, in particular, for $\bar{F}_{CZ,23}$, $\bar{F}_{\frac{\pi}{2},\{1,3\}}$ and $\bar{F}_{\frac{\pi}{2},\{1,2,3\}}$. This may be due to the effect of the crossing point and microwave crosstalk discussed in Appendix H. However, when $\kappa\lesssim 1.5$, all the gates can be implemented with fidelities over 99.99% in the 4-digit precision. Such parasitic capacitances may be experimentally feasible. Therefore, from these results, we expect that the DTC architecture can be extended to multi-qubit systems while keeping its high performance.